write the equation of the line passing through the points (3,-7) with slope -2/3
I set it up like this:
y-(-7)=-2/3(x-3)
y+7=-2/3x-3
y=-2/3x-3-7
y=-2/3x-10
but I got it wrong. It said the b value is -5, How?

Answer by jim_thompson5910(35256)

Your work

y-(-7)=-2/3(x-3)
y+7=-2/3x-3 <--- You made a mistake here
y=-2/3x-3-7
y=-2/3x-10

On the second line, you forgot to distribute the -2/3 to EVERY term inside the parenthesis. You did -2/3 times x correctly, but forgot to multiply -2/3 by -3 to get (-2/3)*(-3) = 6/3 = 2

So line #3 should look like this: y = -2/3x + 2 - 7

and then you combine like terms to get this final answer: y = -2/3x - 5

Despite this error, you did the rest of the steps correctly. So nice work with all that.